Linda L. is in charge of maintaining hospital supplies at General Hospital. During the past year the lead-time when ordering the BX-5 bandage was six weeks. Monthly (four weeks) mean demand was 300 and monthly standard deviation of the demand was 27 units.

  1. What is the reorder point needed to maintain a cycle service level of 99%?
    1. NORMINV(.99,300,27)
    2. NORMINV(.99,300*4/6, 27*4/6)
    3. NORMINV(.99, 300*6/4,27*6/4)

d. None of the above
Answer #1

As the lead time here is 6 weeks which is more than a month period of 4 weeks, therefore the mean and standard deviation of demand here are computed as:

Mean ( 6 week demand ) = (6/4)*300 = 450

Standard deviation ( 6 week demand ) = (6/4)*27 = 40.5

Therefore the reorder point now is computed here as:
= Total demand in 6 weeks which is not exceeded with 99% probability.

In EXCEL this is computed as:
= NORMINV(.99, 300*6/4,27*6/4)

Therefore c is the required expression here.